tmi systems, a company that customizes software for construction cost estimates, repaid a loan obtained 7 years ago at 10% per year simple interest. if the amount that tmi repaid was $180,000. calculate the principal of the loan?
Answer #1

Total amount = Principal + Simple Interest

P+SI = 180,000
SI = PRT/100
P(1+RT/100) = 180,000

P = 18000/(1+0.7)
P = 105882.35